OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). A.) y = 2x + 6 B.) y = -one halfx + 3 C.)y = -one halfx + 6 D.) y = 2x + 3

OpenStudy (anonymous):

What's the slope of the line y = 2x + 2?

OpenStudy (anonymous):

I do not understand

OpenStudy (anonymous):

The slope is by the x variable when the line is written in slope intercept form y = mx+b. In the above equation, 2 is the slope

OpenStudy (anonymous):

To get a perpendicular line, the slope of the new line must be the opposite inverse of the given line. The given line has a slope of 2 The line perpendicular to this will be the opposite inverse or -1/2

OpenStudy (anonymous):

Now...you want the perpendicular line to pass through the point (6,3). Plug everything into this equation: y - y1 = m(x-x1) x1. y1 6 3 y - 3 = -1/2(x - 6) now solve y - 3 = -1/2x + 3 y = -1/2x + 6 This is C in your list

OpenStudy (anonymous):

thank you i get it now

OpenStudy (anonymous):

good luck